• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ellie

    Log entry time 15:57:22 on March 15, 2011

    Entry number 351949

    keyword=Day Shift Summary

    03/15/11 Swing Shift Summary

    On shift:
    SL: Ellie Long
    TO: Todd Averett
    3rd: Tim Holstrom

    · 08:00 Good morning, Hall A! The shift starts in Production with 3.5uA on D target on run 2833.


    · 09:08 Run Coordinator noticed that the Beam Time Accounting didn't seem to be outputting to the HALog. Restarted the BTA GUI.


    · 09:45 OMG A TORNADO!!! RUN!!!! Fleeing the area with the 3rd for safer ground! Target Operator ties himself to a heavy pipe using his leather belt and waits it out. (Actually just a tornado drill)


    · 09:55 Whew! Tornado over! Make it back upstairs safe and sound and the TO is back after his short visit to Oz. (Tornado drill over)


    · 10:00 MCC called to begin Beam Studies. We called for Controlled Access during this time.


    · 15:15 RC asked for Restricted Access, which we've gone into.


    · 16:00 Finished shift with no beam and work being done on the spectrometer magnets. Next shift leader: Be sure to check the kinematics settings with the RC (Alexandre) as I believe they are changing them.





    Production Runs:

     Run #  Current  Target  # of Events  Comments 
    2833 3.5uA 10cm Loop 2 He 1M
    2834 3.5uA 10cm Loop 2 He 1M
    2835 3.5uA 10cm Loop 2 He 1.1M



    Junk Runs:

     Run #  Current  Target  # of Events  Comments 
    2836 ---- 10cm Loop 2 He 70k No beam during run